(32a^9)(4a^-2)/ 128a^2 I am having a real problem with that question because I dont go to a regular school where I am taught a lesson. I go to an alternative school and they just give me the work and expect me to know it. If you could please help me understand it I would really appricate it.
Answer by piyu(4)

your question was (32a^9)(4a^-2)/ 128a^2
it is nothing but (32)(4)(a^9)/(128a^2)(a^2)
=128(a)^4+5 divided by (128a^2+2)
=128(a)^4(a)^5 divided by (128)(a^4)(a^2)
=a^5
tha 128 and a^4 in the numerator and the denominator cancells off and only a^5 remains.